OpenStudy (anonymous):

a sporting good store received an order of 80 baseball caps, of which 12 were green. if 1 of the 80 caps is selected at random, what is the probability it will not be green?

OpenStudy (johnweldon1993):

Alright so all together...we have 80 caps... 12 of which are green....so that is \[\large \frac{12}{80} = \text{green}\] So the probability that they will NOT be green...will be \[\large 1 - \frac{12}{80} = \frac{68}{80}\] And that can be simplified down a bit

OpenStudy (anonymous):

so 85%?

OpenStudy (johnweldon1993):

Correct :)